Determine the general solution if:cos2x-5cosx-2=0
Answer by Cromlix(4381)

Hi there,
cos2x-5cosx-2=0
(cos2x = 2cos^2 - 1)
2cos^2x - 1 - 5cosx - 2
2cos^2x - 5cosx - 3 = 0
(2cosx + 1)(cosx - 3)= 0
cosx - 3 = 0
cosx = 3 (Discount as cosx only
goes up to 1 and down to -1)
2cosx + 1 = 0
2cosx = -1
cosx = -1/2
x = 120 and 240 degrees
x = 2pi/3 or 4pi/3
Hope this helps :-)